OpenStudy (anonymous):

Solve The Equation Beloww ; -30 = q - 10 + 11q

OpenStudy (anonymous):

q= -5/3

OpenStudy (anonymous):

group the like terms so have the variables on one side and the constants on the other so q + 11q = -30 +10 12q = -20 q = -20/12 q= -5/3